Reports 1-1 of 1 Clear search Modify search
VIS (General)
mark.barton - 12:41 Wednesday 20 March 2019 (8468) Print this report
Type B OL Whitening/Dewhitening

Tuesday 3/19 and Wednesday 3/20.

I did an audit of the whitening and dewhitening for the Type B OLs. Yokozawa-san had set up the BS, so I used that as a model.

  • The recipe is to enable one of the three hardware whitening filters on each segment of both LEN and TILT (switch 5 on the BIO Config card) plus any additional gain stages (switch 4 -> 24 dB, 3 -> 12 dB, 2 -> 6 dB, 1 -> 3dB) required to get the SUM up to about 20K-30K counts. (A higher SUM would give better S/N, but a SUM greater than 32K makes it difficult to check for uneven segments by directing the beam to one segment at a time.)
  • Each whitening filter is then compensated for by a filter AntiWTN = FM5 = zpk([10],[1],1,”n”) in the OPLEVINF block.

I found and fixed the following issues:

  • BS LEN had the AntiWTN filter installed in the OPLEVINF blocks but it was not enabled, so I turned it on and did a new safe.snap.
  • SR2, SR3 and SRM did not have the AntiWTN filters in either the LEN OR TILT OPLEVINF blocks, so I copied them from BS (), enabled them, and did a new safe.snap.
  • SR2 did not have the whitening filters enabled, so I turned them on.
  • SRM had the wrong type of whitening chassis (with no whitening card!), S1808108, so the BIO Config cards didn't have any effect. I replaced it with one of the correct type, S1808656.

The final SUM values and switch settings (for now) are:

  • BS LEN: 1600 (beam not on QPD); 5 (WTN), 4 (24 dB), 1 (3 dB).
  • BS TILT: 25700; 5 (WTN), 4 (24 dB).
  • SR2 LEN: 29400; 5 (WTN),  2 (6dB).
  • SR2 TILT: 18600; 5 (WTN), 2 (6dB).
  • SR3 LEN: 17900; 5 (WTN), 4 (24 dB), 1 (3 dB).
  • SR3 TILT: 15400; 5 (WTN), 4 (24 dB), 1 (3 dB).
  • SRM LEN: 18700; 5 (WTN), 2 (6dB).
  • SRM TILT: 14000; 5 (WTN), 2 (6dB).
Search Help